OpenStudy (anonymous):

What are the zeros of the polynomial function f(x) = x3 - 2x2 - 24x?

hero (hero):

f(x) = x^3 - 2x^2 - 24x Set f(x) = 0 0 = x^3 - 2x^2 - 24x Factor out x 0 = x(x^2 - 2x - 24) Factor x^2 - 2x - 24. Then use zero product property to finish solving for x.

OpenStudy (anonymous):

How would I factor it? I'm terrible at this.. :/

hero (hero):

Find 2 numbers that multiply to get 24 but combine to get -2

OpenStudy (anonymous):

-6 and 4?

hero (hero):

Yes correct. So x^2 -2x - 24 = (x - )(x + ) Fill in the blank

OpenStudy (anonymous):

(x-6)(x+4) So the answer would be -6, 0, 4?

hero (hero):

So that you don't make a mistake, write out the steps like this: 0 = x(x - 6)(x + 4) x = 0 x - 6 = 0 x + 4 = 0 x = 0 x = 6 x = -4

OpenStudy (anonymous):

Thank you so much!